OpenStudy (anonymous):

Television sets. What does it mean to refer to a 20 in TV set or 25 in TV set? Such units refer to the diagonal of the screen. A 30 in TV set also has a width of 24 inches. What is its height? What is the height of a 30 in TV? ___ inches

OpenStudy (anonymous):

use the pythagorean theorem, a^2+b^2=c^2 24^2+b^2=30^2 900-2^6(9)=b^2=900-64*9=900-576=324 b=\[\sqrt{324}\] =18in